# DIP-R1: Deep Inspection and Perception with RL Looking Through and Understanding Complex Scenes

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated significant visual understanding capabilities, yet their fine-grained visual perception in complex real-world scenarios, such as densely crowded public areas, remains limited. Inspired by the recent success of reinforcement learning (RL) in both LLMs and MLLMs, in this paper, we explore how RL can enhance visual perception ability of MLLMs. Then we develop a novel RL-based framework, Deep Inspection and Perception with RL (DIP-R1) designed to enhance the visual perception capabilities of MLLMs, by comprehending complex scenes and looking through visual instances closely. DIP-R1 guides MLLMs through detailed inspection of visual scene via three simply designed rule-based reward modelings. First, we adopt a standard reasoning reward encouraging the model to include three step-by-step processes: 1) reasoning for understanding visual scenes, 2) observing for looking through interested but ambiguous regions, and 3) decision-making for predicting answer. Second, a variance-guided looking reward is designed to examine uncertain regions for the second observing process. It explicitly enables the model to inspect ambiguous areas, improving its ability to mitigate perceptual uncertainties. Third, we model a weighted precision-recall accuracy reward enhancing accurate decision-making. We explore its effectiveness across diverse fine-grained object detection data consisting of challenging real-world environments, such as densely crowded scenes. Built upon existing MLLMs, DIP-R1 achieves consistent and significant improvement across various in-domain and out-of-domain scenarios. It also outperforms various existing baseline models and supervised fine-tuning methods. Our findings highlight the substantial potential of integrating RL into MLLMs for enhancing capabilities in complex real-world perception tasks.
1 Introduction
@@
These days, multimodal large language models (MLLMs) have achieved remarkable progress [2, 3, 4, 5, 6, 7, 8], accelerated by advances in large language models (LLMs) [9, 10, 11, 12, 13, 14, 15, 16, 17], large-scale vision-language pretraining [18, 19], instruction tuning [20, 21], and multimodal reasoning [22, 23, 24]. These developments have significantly improved visual understanding capabilities of MLLMs, enabling noticeable performance across a wide range of visual understanding tasks, such as visual question answering (VQA), referring expression comprehension (REC), and image-text retrieval [25, 26]. More recently, following the success of Reinforcement Learning (RL) in improving the reasoning abilities of LLMs, several studies have explored its application to visual domain tasks, demonstrating its effectiveness to enhance visual understanding and generalization with a small amount of data [24, 27, 28, 29, 30, 31].
However, despite these advancements, MLLMs still suffer from perceiving complex real-world scenes, such as public places, where MLLMs can help human decision-making for public safety [32]. Such public scenes usually involve densely populated environments where many instances appear and overlap. For example, Figure 1(a) shows example public scenes and the visual perception (i.e., person detection) results of Qwen2.5-VL-3B-Instruct [1]. As shown in the figure, while it properly perceives each person separately in the relatively easy environment (first example), it fails to capture each individual in crowded scenes, considering them as a single group. To mitigate this problem, MLLMs need to enhance their fine-grained visual perception ability by separating, recognizing, and localizing each instance properly, instead of regarding them as a single entity.
@@
In this paper, we explore the way to enhance MLLMs’ fine-grained visual perception capability, based on recent findings and open-sources about RL [33, 24, 28, 27]. By utilizing the advantages of RL, we encourage MLLMs to perceive visual scene semantics and each object separately in challenging real-world environments. Specifically, we aim to explicitly guide MLLM to articulate its step-by-step processes. To this end, we employ three simple rule-based reward modelings. First, we employ a standard format reward for MLLM to include three step-by-step processes for each reasoning, observing, and decision-making. MLLM looks through scenes and explains visual semantics especially in the scene level, such as, scene places, instance arrangements, and so on. Second, we design a variance-guided looking reward for the observing process. Here, the variance is estimated to represent how much MLLM is uncertain its response, especially about region prediction. Due to this variance-guided looking process, the observing process is reinforced to include uncertain regions (which have high variances). Therefore, it enables MLLM to densely inspect uncertain region candidates which are confusing for being recognized properly. Third, we incorporate a weighted precision-recall reward for decision-making process to minimize missing instances, predict correct region candidates, and give accurate answer. By integrating these components, we develop a Deep Inspection and Perception RL framework, called DIP-R1, which can look through and comprehend visual scenes within challenging real-world environments.
@@
We investigate the visual understanding ability in complex scenes with MLLMs’ object detection performance in wild scenarios including densely crowded scenes, driving street areas, and aerial safety monitoring views, and so on. Extensive experimental results demonstrate that the proposed DIP-R1 framework is significantly effective in boosting MLLMs’ visual perception capability in such challenging environments. DIP-R1 framework achieves noticeable improvements on four real-world object detection datasets, CrowdHuman [34], CityPersons [35], WiderPedestrian [36], and UAVDT [37]. Figure 1(b) shows detection performance on CrowdHuman and comparison with other baseline models and supervised fine-tuning (SFT) method on same Qwen2.5-VL baseline. In this work, we explore how RL is effective in improving visual perception capability of MLLMs in complex scenes. The contribution of our work can be summarized as follows:
@@
  • We incorporate three simple yet effective rule-based reward modelings: 1) a format reward to enforce reasoning, observing, and decision-making steps, 2) a variance-guided looking reward to promote inspection of uncertain regions, and 3) a weighted precision-recall reward to perform robust instance-level recognition by securing balanced precision and recall.
  • We demonstrate the effectiveness of DIP-R1 through comprehensive experiments on four challenging real-world detection benchmarks, achieving large performance gain over a variety of existing baseline models including SFT.
2 Related Works
2.1 Multimodal Large Language Models (MLLMs)
@@
2.2 Reinforcement Learning (RL)
RL has been considered a strong machine learning method, enabling agents to learn how to interact with environments and make decisions based on predefined reward policy [54]. More recently, the emergence of LLMs and powerful open-sourcing foundations, such as OpenAI-o1 [55] and VLM-R1 [27], has accelerated significant progress in applying RL for aligning language models with human preferences. Pioneering approaches include Reinforcement Learning with Human Feedback (RLHF) [56, 57], Proximal Policy Optimization (PPO) [58], Direct Preference Optimization (DPO) [59], and Kahneman-Tversky Optimization (KTO) [60]. Recent works (e.g., ReST-MCTS [61] and DeepSeek-R1 [33]) have demonstrated that simple rule-based or outcome-level reward signals can effectively improve the reasoning abilities of LLMs. In particular, DeepSeek-R1 [33] employ Group Relative Policy Optimization (GRPO) [62], which aggregates and compares group-level and token-level preferences. Following these advances, RL has been increasingly extended to MLLMs. While early efforts mainly focused on mitigating hallucinations and improving alignment through preference feedback, more recent works enhance multimodal reasoning and understanding via RL-based fine-tuning. For example, R1-OneVision [63] aims to bridge the gap between visual perception and reasoning by integrating Chain-of-Thought (CoT) approach with GRPO. R1-VL [28] presents StepGRPO, which improves structured reasoning by providing step-wise accuracy and validity rewards. Visual-RFT [29] moves beyond mathematical reasoning to visual understanding of MLLMs, leveraging a rule-based Intersection-of-Union (IoU) reward function with GRPO. VLM-R1 [27] also incorporates GRPO into visual domain to enhance MLLMs’ generalization and understanding of open-world instances. Building on this line of research, our work explores how RL can further improve the visual perception abilities of MLLMs, particularly in complex and challenging environments.
3 Preliminaries
3.1 Group Relative Policy Optimization (GRPO)
@@
3.2 Uncertainty Estimation in Neural Networks
In deep neural networks (DNNs), various studies have investigated how to measure the variance of model predictions [64, 65, 66]. For example, epistemic uncertainty has been introduced to estimate how much the deep learning model is uncertain about its prediction because of limited training data or imperfect learning [64]. This kind of uncertainty, which is also referred to as predictive uncertainty, estimates the uncertainty in the model parameters [65]. A common approach to measure this uncertainty is by Monte-Carlo (MC) dropout approximation [66], where the model’s outputs are sampled multiple times. Specifically, while sampling $T$ predictions, the predictive mean is measured by $\mathbb{E}[y]\approx\frac{1}{T}\sum_{t=1}^{T}\hat{y}_{t}$, where $\hat{y}_{t}$ is the model’s prediction probability at the $t$-th sampling. The corresponding predictive variance–epistemic uncertainty–is calculated by $\text{Var}[y]\approx\frac{1}{T}\sum_{t=1}^{T}\hat{y}_{t}^{2}-\left(\mathbb{E}[y]\right)^{2}$. Given that GRPO is also a sampling-based algorithm which generates $G$ group responses for policy optimization, a question naturally arises: “Can we also estimate the model’s variance during GRPO?”. Inspired by this conceptual similarity, in this work, we explore whether MLLMs can assess the variance across group responses during GRPO, and whether this variance can be effectively used as a reward signal to improve visual understanding.
4 DIP-R1 Framework
In this section, we describe the details about our proposed framework, Deep Inspection and Perception RL framework (DIP-R1). Figure 2 illustrates the overall architecture, which incorporates GRPO with three simple rule-based reward modelings: 1) think-look-answer format reward which encourages step-by-step reasoning, observing, and decision-making processes, 2) variance-guided look reward which helps look through uncertain regions, and 3) Weighted Precision-Recall Accuracy Reward which considers both precision and recall to predict accurate answer. Our default instruction template and example output format are shown below. The more details about our reward modelings are elaborated in the following subsections.
[…]
@@ 4 DIP-R1 Framework
4.1 Think-Look-Answer Format Reward
We design the think-look-answer format reward to encourage MLLM to infer the answer through three step-by-step processes: scene-level reasoning, region-level observing, and decision-making. Based on the classic format reward for <think> … </think><answer> … </answer> [33, 67, 27, 68], we add <look> … </look> between them. Therefore, MLLM learns to generate responses following three step-by-step processes: reasoning, observing, and decision-making. The Think-Look-Answer reward assigns 1 if the response includes all three components in the correct order, and 0 otherwise:
@@ 4.1 Think-Look-Answer Format Reward
4.2 Variance-guided Look Reward
@@
$r_{\mathbf{b}}=\begin{cases}\alpha\text{Var}[\mathbf{k_{b}}],&\text{if }\text{Var}[\mathbf{b}]\neq 0,\\ -1/\alpha,&\text{otherwise},\end{cases}\qquad\mathrm{R}_{\text{look}}=\frac{1}{N_{L}}\sum_{i=1}^{N_{L}}r_{\mathbf{b}_{i}},$ (5)
where $N_{L}$ is the number of region boxes observed in the observing process, and $\alpha$ is a normalization factor (dependent on $G$) to scale the reward to [0, 1] when $\text{Var}[\mathbf{k_{b}}]\neq 0$. Otherwise, a small penalty is applied when certain regions are included (i.e., $\text{Var}[\mathbf{k_{b}}]=0$). In summary, the more uncertain regions are involved in the observing process, the higher reward is be acquired. As a result, due to this variance-guided look reward, the observing process learns to inspect ambiguous regions in complex scenes that MLLM is confused. The pseudo-code for variance-guided look reward is described in Appendix A.
4.3 Precision-Recall Accuracy Reward
@@
$\mathrm{R}_{\text{accuracy}}=w_{r}\cdot\boldsymbol{P}(\mathcal{B}_{pred},\mathcal{B}_{gt}),\qquad w_{r}=1+\text{bin}\left[\boldsymbol{R}(\mathcal{B}_{pred},\mathcal{B}_{gt})\right].$ (6)
$\boldsymbol{P}(\cdot)$ and $\boldsymbol{R}(\cdot)$ are the functions calculating precision and recall metrics, respectively. $\text{bin}\left[\boldsymbol{R}(\mathcal{B}_{pred},\mathcal{B}_{gt})\right]$ discretizes a recall value into one of 100 equal intervals over the range [0, 1], so that $w_{r}$ denotes a weight which is linearly increasing depending on recall. Since a simple precision reward modeling considers precision only, so that it fails to decrease false negatives. Different from the simple precision reward, this reward design takes into account recall value, so that it becomes large when both precision and recall are high. As a result, as the model’s answer is more accurate (i.e., high precision) and covers as many as relevant regions (i.e., high recall), this reward modeling can assign large reward to the policy. Owing to this reward, our DIP-R1 framework generates more accurate answers even within crowded scenes where many objects appear and overlapped.
5 Experiments
5.1 Datasets and Implementation Details
Table 1: Performance comparison with baseline models on in-domain CrowdHuman [34] and out-of-domain CityPersons [35], WiderPedestrian [36], and UAVDT [37]. The best results are highlighted in bold and the runner-up results are underlined.
@@ 5.1 Datasets and Implementation Details
CrowdHuman (In-domain data) CityPersons (Out-of-domain data)
@@
WiderPedestrian (Out-of-domain data) UAVDT (Out-of-domain data)
@@
To construct our framework, we adopt Qwen2.5-VL-Instruct (3B and 7B) [1] as our base model–one of the pioneering and open-sourcing MLLMs. We compare the performance of our DIP-R1 framework with several existing methods, including Ferret (7B) [40], InternVL3 (8B) [70], DeepSeek-VL2-Small (16B) [23], and Groma (7B) [53]. For training, we use 4 NVIDIA A100 80GB GPUs. The details of our GRPO training configuration are: number of generations $G$ is 8, batch size per device is 4, gradient accumulation steps are 4, the number of training epoch is 1, KL divergence coefficient $\beta$ is 0.04, maximum output length is 2048 tokens, and the learning rate is 1e-6. We report the best performance of our framework and SFT method within an epoch for the main comparison in Table 1. For evaluating baseline models, we follow their own default prompts, and further details are described in Appendix C. DIP-R1 framework is built upon well-established RL foundations [27].
5.2 Experimental Results
5.2.1 Comparisons with Baselines
In this section, we compare the performance of our DIP-R1 framework with various baseline models on four object detection datasets, including in-domain data: CrowdHuman [34] and three out-of-domain data: CityPersons [35], WiderPedestrian [36], and UAVDT [37]. Table 1 shows the performance comparison with the other baseline models: Ferret (7B) [40], Qwen2.5-VL-Instruct (3B, 7B) [1], InternVL3 (8B) [70], DeepSeek-VL2-Small (16B) [23], and Groma (7B) [53]. On CrowdHuman, our framework shows remarkable performance gains over existing baselines, including DeepSeek-VL2-Small (16B) [23] and InternVL3 (8B) [70], which are among the most recently released state-of-the-art MLLMs. Notably, even compared to Groma [53]–a grounding-specialized MLLM which adopts additional region proposal networks, our framework achieves superior performance. On the out-of-domain data, DIP-R1 framework consistently enhances visual perception capabilities across diverse, challenging environments.
Figure 3: The visualization results of Qwen2.5-VL (3B) [1], SFT on Qwen2.5-VL (3B), DeepSeek-VL2-Small (16B) [23], and DIP-R1 (3B), along with ground-truth regions. As shown in the figure, the existing baseline models easily fail to perceive each individual and tend to miss or consider them as a single entity. On the other hand, our method separates and recognizes each of them properly.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated significant visual understanding capabilities, yet their fine-grained visual perception in complex real-world scenarios, such as densely crowded public areas, remains limited. Inspired by the recent success of reinforcement learning (RL) in both LLMs and MLLMs, in this paper, we explore how RL can enhance visual perception ability of MLLMs. Then we develop a novel RL-based framework, Deep Inspection and Perception with RL (DIP-R1) designed to enhance the visual perception capabilities of MLLMs, by comprehending complex scenes and looking through visual instances closely. DIP-R1 guides MLLMs through detailed inspection of visual scene via three simply designed rule-based reward modelings. First, we adopt a standard reasoning reward encouraging the model to include three-step reasoning process: 1) comprehending entire visual scene, 2) observing for looking through interested but ambiguous regions, and 3) decision-making for predicting answer. Second, a variance-guided looking reward is designed to encourage MLLM to examine uncertain regions during the second observation process, guiding it to inspect ambiguous areas and thereby mitigate perceptual uncertainty. This reward further promotes variance-driven visual exploration, enabling the MLLM to reason about region-level uncertainty and explicitly indicate interpretable uncertain regions. Third, we model a weighted precision-recall accuracy reward enhancing accurate decision-making. We explore its effectiveness across diverse fine-grained object detection data consisting of challenging real-world environments, such as densely crowded scenes. Built upon existing MLLMs, DIP-R1 achieves consistent and significant improvement across various in-domain and out-of-domain scenarios. It also outperforms various existing baseline models and supervised fine-tuning methods. Our findings highlight the substantial potential of integrating RL into MLLMs for enhancing capabilities in complex real-world perception tasks.
I Introduction
These days, multimodal large language models (MLLMs) have achieved remarkable progress [1, 2, 3, 4, 5, 6, 7, 8, 9], accelerated by advances in large language models (LLMs) [10, 11, 12, 13, 14, 15, 16, 17, 18], large-scale vision-language pretraining [19, 20, 21], instruction tuning [22, 23], and multimodal reasoning [24, 25, 26, 27, 28]. These developments have significantly improved visual understanding capabilities of MLLMs, enabling noticeable performance across a wide range of visual understanding tasks, such as visual question answering (VQA), referring expression comprehension (REC), and image-text retrieval [29, 30, 31]. More recently, following the success of Reinforcement Learning (RL) in improving the reasoning abilities of LLMs, several studies have explored its application to visual domain tasks, demonstrating its effectiveness to enhance visual understanding and generalization with a small amount of data [26, 32, 33, 34, 35, 36]. However, despite these advancements, MLLMs still suffer from perceiving complex real-world scenes, such as public places, where MLLMs can help human decision-making for public safety [37]. Such public scenes usually involve densely populated environments where many instances appear and overlap. For example, Fig. 1(a) shows example public scenes and the visual perception (i.e., person detection) results of Qwen2.5-VL-3B-Instruct [38]. As shown in the figure, while it properly perceives each person separately in the relatively easy environment (described in the first top-left example), it fails to capture each individual in crowded scenes, considering them as a single group. To mitigate this problem, MLLMs need to enhance their fine-grained visual perception ability by separating, recognizing, and localizing each instance properly, instead of regarding them as a single entity.
@@ I Introduction
In this paper, we explore the way to enhance MLLMs’ fine-grained visual perception capability, based on recent findings and open-sources about RL [39, 26, 33, 32]. By utilizing the advantages of RL, we encourage MLLMs to perceive visual scene semantics and each object separately in challenging real-world environments. Specifically, we aim to explicitly guide MLLM to articulate its step-by-step processes. To this end, we employ three simple rule-based reward modelings. First, we employ a standard format reward for MLLM to include three step-by-step processes for each reasoning, observing, and decision-making. MLLM looks through scenes and explains visual semantics especially in the scene level, such as, scene places, instance arrangements, and so on. Second, we design a variance-guided looking reward for the observing process. Here, the variance is estimated to represent how much MLLM is uncertain its response, especially about region prediction. Due to this variance-guided looking process, the observing process is reinforced to include uncertain regions (which have high variances). Therefore, it enables MLLM to densely inspect uncertain region candidates which are confusing for being recognized properly. In other words, this variance-driven visual exploration guides MLLM to reason over region-level uncertainty. Third, we incorporate a weighted precision-recall reward for decision-making process to minimize missing instances, predict correct region candidates, and give accurate answer. By integrating these components, we develop Deep Inspection and Perception RL framework, named DIP-R1, which can look through and comprehend visual scenes within challenging real-world environments.
@@
We investigate the visual understanding ability in complex scenes with MLLMs’ object detection performance in wild scenarios including densely crowded scenes, driving street areas, and aerial safety monitoring views, and so on. Extensive experimental results demonstrate that the proposed DIP-R1 framework is significantly effective in boosting MLLMs’ visual perception capability in such challenging environments. DIP-R1 framework achieves noticeable improvements on four real-world object detection datasets, CrowdHuman [40], CityPersons [41], WiderPedestrian [42], and UAVDT [43]. Fig. 1(b) shows detection performance on CrowdHuman and comparison with other baseline models and supervised fine-tuning (SFT) method on same Qwen2.5-VL baseline. In this work, we explore how RL is effective in improving visual perception capability of MLLMs in complex scenes. The contribution of our work can be summarized as follows:
@@
  • We incorporate three simple yet effective rule-based reward modelings: 1) a format reward to enforce reasoning, observing, and decision-making: three-step reasoning process, 2) a variance-guided looking reward to promote inspection of uncertain regions, and 3) a weighted precision-recall reward to perform robust instance-level recognition by securing balanced precision and recall.
  • The explicit variance-driven visual exploration is incorporated in the reasoning process of MLLM, so that it enables MLLM to inspect uncertain regions and reason over region-level uncertainty by itself.
  • We demonstrate the effectiveness of DIP-R1 through comprehensive experiments on four challenging real-world detection benchmarks, achieving large performance gain over existing baseline models including SFT.
II Related Work
II-A Multimodal Large Language Models (MLLMs)
@@
II-B Reinforcement Learning (RL)
RL has been considered a strong machine learning method, enabling agents to learn how to interact with environments and make decisions based on predefined reward policy [60]. More recently, the emergence of LLMs and powerful open-sourcing foundations, such as OpenAI-o1 [61] and VLM-R1 [32], has accelerated significant progress in applying RL for aligning language models with human preferences. Pioneering approaches include Reinforcement Learning with Human Feedback (RLHF) [62, 63], Proximal Policy Optimization (PPO) [64], Direct Preference Optimization (DPO) [65], and Kahneman-Tversky Optimization (KTO) [66]. Recent works (e.g., ReST-MCTS [67] and DeepSeek-R1 [39]) have demonstrated that simple rule-based or outcome-level reward signals can effectively improve the reasoning abilities of LLMs. In particular, DeepSeek-R1 [39] employ Group Relative Policy Optimization (GRPO) [68], which aggregates and compares group-level and token-level preferences. Following these advances, RL has been increasingly extended to MLLMs. While early efforts mainly focused on mitigating hallucinations and improving alignment through preference feedback, more recent works enhance multimodal reasoning and understanding via RL-based fine-tuning. For example, R1-OneVision [69] aims to bridge the gap between visual perception and reasoning by integrating Chain-of-Thought (CoT) approach with GRPO. R1-VL [33] presents StepGRPO, which improves structured reasoning by providing step-wise accuracy and validity rewards. Visual-RFT [34] moves beyond mathematical reasoning to visual understanding of MLLMs, leveraging a rule-based Intersection-of-Union (IoU) reward function with GRPO. VLM-R1 [32] also incorporates GRPO into visual domain to enhance MLLMs’ generalization and understanding of open-world instances.
Building on this line of research, our work explores how RL can further improve the visual perception abilities of MLLMs, particularly in complex and challenging environments. Different from existing RL-based works, our framework incorporate additional reasoning process (i.e., observing process)–a variance-driven visual exploration helping MLLM inspect and reason over uncertain regions by itself.
III Preliminary
III-A Group Relative Policy Optimization (GRPO)
@@
III-B Uncertainty Estimation in Neural Networks
In deep neural networks (DNNs), various studies have investigated how to measure the variance of model predictions [70, 71, 72]. For example, epistemic uncertainty has been introduced to estimate how much the deep learning model is uncertain about its prediction because of limited training data or imperfect learning [70]. This kind of uncertainty, which is also referred to as predictive uncertainty, estimates the uncertainty in the model parameters [71]. A common approach to measure this uncertainty is by Monte-Carlo (MC) dropout approximation [72], where the model’s outputs are sampled multiple times. Specifically, while sampling $T$ predictions, the predictive mean is measured by:
$\mathbb{E}[y]\approx\frac{1}{T}\sum_{t=1}^{T}\hat{y}_{t},$ (2)
where $\hat{y}_{t}$ is the model’s prediction probability at the $t$-th sampling. The corresponding predictive variance–epistemic uncertainty–is calculated by:
$\text{Var}[y]\approx\frac{1}{T}\sum_{t=1}^{T}\hat{y}_{t}^{2}-\left(\mathbb{E}[y]\right)^{2}.$ (3)
Given that GRPO is also a sampling-based algorithm which generates $G$ group responses for policy optimization, a question naturally arises: “Can we also estimate the model’s variance during GRPO?”. Inspired by this conceptual similarity, in this work, we explore whether MLLMs can assess the variance across group responses during GRPO, and whether this variance can be effectively used as a reward signal to improve visual understanding.
IV Proposed Framework: DIP-R1
In this section, we describe the details about our proposed framework, Deep Inspection and Perception RL framework (DIP-R1). Fig. 2 illustrates the overall architecture, which incorporates GRPO with three simple rule-based reward modelings: 1) think-look-answer format reward which encourages step-by-step reasoning, observing, and decision-making processes, 2) variance-guided look reward which helps look through uncertain regions, and 3) Weighted Precision-Recall Accuracy Reward which considers both precision and recall to predict accurate answer. Our default instruction template and example output format are shown below. The more details about our reward modelings are elaborated in the following subsections.
@@ IV Proposed Framework: DIP-R1
Example Output Format <think> First, I’$l$ inspect the image to identify each person present. The group consists of around 20 individuals standing closely and posing for a photo. Since the task is to identify individual persons in the crowd, $\cdots$ </think> <look> [{“bbox_2d”: [124, 423, 183, 693], “label”: “person”}, {“bbox_2d”: $\cdots$] </look> <answer> [{“bbox_2d”: [258, 451, 343, 845], “label”: “person”}, {“bbox_2d”: $\cdots$] </answer>
IV-A Think-Look-Answer Format Reward
We design the think-look-answer format reward to encourage MLLM to infer the answer through three step-by-step processes: scene-level reasoning, region-level observing, and decision-making. Based on the conventional format reward for <think> ... </think><answer> ... </answer>, we add <look> ... </look> between them. Therefore, MLLM learns to generate responses following three step-by-step processes: reasoning, observing, and decision-making. The Think-Look-Answer reward assigns 1 if the response includes all three components in the correct order, and 0 otherwise:
@@ IV-A Think-Look-Answer Format Reward
IV-B Variance-guided Look Reward
@@
$r_{\mathbf{b}}=\begin{cases}\alpha\text{Var}[\mathbf{k_{b}}],&\text{if }\text{Var}[\mathbf{b}]\neq 0,\\ -1/\alpha,&\text{otherwise},\end{cases}$ (7)
$\mathrm{R}_{\text{look}}=\frac{1}{N_{L}}\sum_{i=1}^{N_{L}}r_{\mathbf{b}_{i}},$ (8)
where $N_{L}$ is the number of region boxes observed in the observing process, and $\alpha$ is a normalization factor (dependent on $G$) to scale the reward to [0, 1] when $\text{Var}[\mathbf{k_{b}}]\neq 0$. Otherwise, a small penalty is applied when certain regions are included (i.e., $\text{Var}[\mathbf{k_{b}}]=0$). In summary, the more uncertain regions are involved in the observing process, the higher reward is be acquired. As a result, due to this variance-guided look reward, the observing process learns to inspect ambiguous regions in complex scenes that MLLM is confused. Algorithm 1 is the pseudo-code elaborating how to compute the variance-guided look reward in details.
Algorithm 1 The way to Compute Variance-guided Look Reward
2:Initialize set of all boxes $\mathcal{B}\leftarrow\bigcup_{i=1}^{G}\mathcal{B}_{i}$ $\triangleright$ All predicted boxes from $G$ responses
3:for each box $\mathbf{b}\in\mathcal{B}$ do
4:  $f_{b}\leftarrow 0$
5:  for $i=1,\dots,G$ do
7:     $f_{b}\leftarrow f_{b}+1$ $\triangleright$ Count the frequency of same box across generations
8:     continuecontinue
9:   end if
10:  end for
11:  $f_{b}\leftarrow f_{b}/G$
12:  $\text{Var}[\mathbf{k}_{\mathbf{b}}]\leftarrow f_{b}(1-f_{b})$ $\triangleright$ Compute variance
13:  if $\text{Var}[\mathbf{k}_{\mathbf{b}}]>0$ then
14:   $r_{\mathbf{b}}\leftarrow\alpha\cdot\text{Var}[\mathbf{k}_{\mathbf{b}}]$ $\triangleright$ Compute box reward based on the variance
15:  else
16:   $r_{\mathbf{b}}\leftarrow-1/\alpha$ $\triangleright$ Penalty for certain boxes
17:  end if
18:end for
[…]
21:  Let $\mathcal{B}^{l}_{i}=\{\mathbf{b}_{1},\dots,\mathbf{b}_{N_{L}}\}$
22:  $\mathrm{R}_{i}\leftarrow\frac{1}{N_{L}}\sum_{j=1}^{N_{L}}r_{\mathbf{b}_{j}}$ $\triangleright$ Compute Look Reward for each response
23:  $\mathrm{R}_{\text{look}}\text{.append}(\mathrm{R}^{i})$
24:end for
25:return $\mathrm{R}_{\text{look}}$
IV-C Precision-Recall Accuracy Reward
@@
$\mathrm{R}_{\text{accuracy}}=w_{r}\cdot\boldsymbol{P}(\mathcal{B}_{pred},\mathcal{B}_{gt}),$ (9)
$w_{r}=1+\text{bin}\left[\boldsymbol{R}(\mathcal{B}_{pred},\mathcal{B}_{gt})\right],$ (10)
where $\boldsymbol{P}(\cdot)$ and $\boldsymbol{R}(\cdot)$ are the functions calculating precision and recall metrics, respectively. $\text{bin}\left[\boldsymbol{R}(\mathcal{B}_{pred},\mathcal{B}_{gt})\right]$ discretizes a recall value into one of 100 equal intervals over the range [0, 1], so that $w_{r}$ denotes a weight which is linearly increasing depending on recall. Since a simple precision reward modeling considers precision only, so that it fails to decrease false negatives. Different from the simple precision reward, this reward design takes into account recall value, so that it becomes large when both precision and recall are high. As a result, as the model’s answer is more accurate (i.e., high precision) and covers as many as relevant regions (i.e., high recall), this reward modeling can assign large reward to the policy. Owing to this reward, our DIP-R1 framework generates more accurate answers even within crowded scenes where many objects appear and overlapped.
TABLE I: Performance comparison with baseline models on in-domain (ID) CrowdHuman [40] and out-of-domain (OOD) CityPersons [41], WiderPedestrian [42], and UAVDT [43]. The best results are highlighted in bold and the runner-up results are underlined.
@@
CrowdHuman (ID) CityPersons (OOD)
@@
Model WiderPedestrian (OOD) UAVDT (OOD)
@@
V Experiment
V-A Datasets
V-B Implementation Details
To construct our framework, we adopt Qwen2.5-VL-Instruct (3B and 7B) [38] as our base model–one of the pioneering and open-sourcing MLLMs. We compare the performance of our DIP-R1 framework with several existing methods, including Ferret (7B) [46], InternVL3 (8B) [74], DeepSeek-VL2-Small (16B) [25], and Groma (7B) [59]. For training, we use 4 NVIDIA A100 80GB GPUs. The details of our GRPO training configuration are: number of generations $G$ is 8, batch size per device is 4, gradient accumulation steps are 4, the number of training epoch is 1, KL divergence coefficient $\beta$ is 0.04, maximum output length is 2048 tokens, and the learning rate is 1e-6. The rollout number is $4$ devices × $\times\,4$ batch sizes/device × $\times\,4$ accumulation steps × $\times\,8$ generations = 512=\,512. We report the best performance of our framework and SFT method within an epoch for the main comparison in TABLE I. DIP-R1 framework is built upon well-established RL foundation [32].
Figure 3: The visualization results of Qwen2.5-VL-Instruct (3B), SFT on Qwen2.5-VL-Instruct (3B), DeepSeek-VL2-Small (16B), and DIP-R1 (3B), along with ground-truth regions. As shown in the figure, the existing baseline models easily fail to perceive each individual and tend to miss or consider them as a single entity. On the other hand, our method separates and recognizes each of them properly.